Previous

Calendar

November

2004

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

31

JD305(Snow)
Daylight savings time ends.

1

JD306(Snow/Cold )
~23:00 MST, data system crashed.

2

JD307
Site Visit (SpB): Details , Photos (7)

(hydra down)

~15:00 MST, restarted data system
(ops19, w/ only one edgeport, nwt041102.220251)

3

JD308
Site Visit (SpB, ML, MW): Details , Photos (5)

at 14:00 MST, hydra back up

at 14:00 MST, back to ops18, two edgeports, nwt041103.205212

4

JD309

5

JD310
Site Visit (SpB, JH, JM): Details

at ~12:00 MST, LI-COR LI-7500 at 2.5m was removed to send back to LI-COR for solar radiation fix/repair

East/West Robot flasks swapped (JM)

at 10:00 MST, dry ice test w/ hydra

6

JD311(Warm/Sunny)
at ~9:00 MST, Added 2.4m air inlet to hydra data.

7

JD312
Site Visit (SpB): Details , Photos (7)

Hydra changed to suck, then sample mode. Cal1 stopped, Cal2 started.

Connected the borrowed GMP-343 to UC tower (diff, chan3)

8

JD313

9

JD314
Site Visit (SpB): Details , Photos (17)

Added E/W/N/S/Center inlets to Hydra.

Changed GMP-343 to SE measurement (SE, chan5, TWOSON2.CSI)

10

JD315(Snow )

11

JD316

12

JD317
Site Visit (SpB, ML, HP, KC): Details , Photos (12)

Using new zero-air cylinder for Hydra ref gas (1800 psi).

13

JD318

14

JD319

15

JD320

16

JD321

17

JD322
Robots Start

18

JD323(Sunny/Calm )
Site Visit (SpB): Details , Photos (42)

Turned on fan in the LI-7000 Box, started using c4 (500 ppm).

disconnected KZ PSP.

19

JD324

20

JD325

21

JD326

22

JD327
c4 (500 ppm) stops flowing at time=327.2083 MST

23

JD328

24

JD329
Robots End

25

JD330
TURKEY DAY

26

JD331

27

JD332(Snow)

28

JD333(Snow)

29

JD334

30

JD335
Site Visit (SpB): Details , Photos (53)

c4 (500 ppm) tubing was severed.

1

JD336

2

JD337

3

JD338
Site Visit (SpB): Details , Photos (30)

Replaced Pump for li-6262 (old one died).

connected 748ppm cal gas to c3.

4

JD339